### v1

# Autoregressive models for panel data causal inference with application to state-level opioid policies

Abstract
Motivated by the study of state opioid policies, we propose a novel approach that uses autoregressive models for causal effect estimation in settings with panel data and staggered treatment adoption. Specifically, we seek to estimate of the impact of key opioid-related policies by quantifying the effects of must access prescription drug monitoring programs (PDMPs), naloxone access laws (NALs), and medical marijuana laws on opioid prescribing. Existing methods, such as differences-in-differences and synthetic controls, are challenging to apply in these types of dynamic policy landscapes where multiple policies are implemented over time and sample sizes are small. Autoregressive models are an alternative strategy that have been used to estimate policy effects in similar settings, but until this paper have lacked formal justification. We outline a set of assumptions that tie these models to causal effects, and we study biases of estimates based on this approach when key causal assumptions are violated. In a set of simulation studies that mirror the structure of our application, we also show that our proposed estimators frequently outperform existing estimators. In short, we justify the use of autoregressive models to provide robust evidence on the effectiveness of four state policies in combating the opioid crisis.
Autoregressive models,
Causal inference,
Panel data,
keywords:
[…]
, , , , , , , , , , , and
1 Introduction
The U.S. continues to confront an opioid epidemic (Volkow and Collins, 2017). Between 1999 and 2019, nearly 500,000 people died from an overdose involving opioids (Centers for Disease Control and Prevention (2020a)). The number of fatal overdoses increased nearly 30% from 2019 to 2020, reflecting the spread of fentanyl and greater opioid misuse associated with the increased stress, social isolation, and job loss stemming from the COVID-19 pandemic (Centers for Disease Control and Prevention, 2020b; Patel, Walter and Li, 2021). Provisional 2024 data suggest that overdose deaths will set yet another record (Ahmad et al., 2024). In response, states continue to implement an array of policies meant to combat the opioid-crisis, producing a policy landscape that is both complex and dynamic (Davis and Carr, 2017; Davis and Lieberman, 2021; Davis and Samuels, 2020, 2021; Haffajee, Bohnert and Lagisetty, 2018; Burris et al., 2017; Smart, Haffajee and Davis, 2022; of Emergency Physicians, n.d.; Pacula and Stein, 2020; Andraka-Christou et al., 2022a, b; Frank, Humphreys and Pollack, 2021). Unfortunately, it is unclear whether any of these policies have been effective, as existing research has produced divergent effect estimates of varying credibility. Producing credible effect estimates of these policies is crucial as policymakers choose strategies to address this crisis. In this paper, we therefore quantify the effects of four policies on opioid prescriptions: must access prescription drug monitoring programs (PDMPs), two types of naloxone access laws (NALs), and medical marijuana laws.
Unfortunately, obtaining credible estimates of the causal impacts of opioid-related policies is challenging. The statistical literature on estimating causal effects in panel data settings is extensive, and many existing methods are – broadly speaking – implementations of difference-in-differences (DiD) designs (Ashenfelter and Card, 1984; Angrist and Pischke, 2008) or variations of synthetic control methods (Abadie, Diamond and Hainmueller, 2010, 2015). These include extensions for effect estimation in the staggered adoption of policy setting that we consider (Callaway and Sant’Anna, 2021; Goodman-Bacon, 2021; Ben-Michael, Feller and Rothstein, 2021, 2022). Despite this, these methods are not well suited to the policy landscape seen in the opioid policy evaluation, where multiple policies are simultaneously implemented to address the opioid crisis, and disentangling these effects is important (Griffin et al., 2022). While recent approaches have tried to extend synthetic controls or DiD to multiple policies (Fricke, 2017; Agarwal, Shah and Shen, 2020; Roller and Steinberg, 2023), these are only applicable in certain situations, including when only one of the multiple treatments is of interest. Additionally, even in settings with a single policy, existing estimators have limited statistical power (Griffin et al., 2021, 2023).
To address these shortcomings, Schell, Griffin and Morral (2018) and Griffin et al. (2021) proposed instead using autoregressive models for causal effect estimation. These models allow for more precise estimates of the policies effectiveness while also allowing for estimating effects of multiple policies jointly. A number of simulation studies have shown that these models perform well relative to state-of-the-art implementations of difference-in-differences and synthetic controls when evaluating policy effects (Schell, Griffin and Morral (2018)). Other studies have also shown good performance under different confounding mechanisms, the presence of co-occurring policies, and treatment effect heterogeneity (Griffin et al., 2022, 2023; Griswold et al., 2023). Unfortunately, the challenge with standard autoregressive models is that they control for lags of the dependent variable: in an analysis of policy effects, this will induce “collider bias” due to the fact that the lagged outcomes are likely affected by prior treatments (Achen, 2000; Elwert and Winship, 2014; Pearl, 2009; Nickell, 1981). So-called “debiased autoregressive models” address this problem by removing policy effects from prior outcomes in the estimation process. In this paper we formalize the assumptions behind this method that enable causal identification.
Our primary methodological contribution is therefore to propose a novel causal identification result that justifies the use of debiased autoregressive models for causal effect estimation. Additionally, we study biases of estimates based on this approach when the causal assumptions are violated, particularly with respect to treatment effect heterogeneity. We find that the different estimators in this setting present a standard bias-variance trade-off and that debiased autoregressive models make a very favorable trade-off. To be precise, we show that bias is typically low for these estimators, and they can have significantly lower variance than existing estimators. This is critically important for the analysis of state opioid policy data, where both sample sizes and policy effect sizes may be small and power becomes a pressing issue. Additionally, we formalize the assumptions required to use these models to estimate effects of multiple policies simultaneously, a challenging and relevant problem in opioid policy research. While we primarily seek to demonstrate that this method is useful to quantify the impact of opioid policies, we believe that these models are useful in more general settings.
Our paper proceeds as follows. In Section 2, we outline a causal framework for using debiased autoregressive models to estimate causal effects of a single policy. We also consider violations of these assumptions, particularly the potentially biasing role of effect heterogeneity. In Section 3, we relax the assumptions and show various extensions we can make to the identification result, including estimating the effects of multiple policies jointly. In Section 4, we discuss connections between our approach and other commonly used approaches, including difference-in-differences, synthetic controls, and the methods based on sequential randomization Robins, Hernan and Brumback (2000). In Section 5, we provide a set of simulation studies that further supports the use of our method for our application. In Section 6, we estimate the effects of state policies on opioid prescriptions and compare our estimates against other existing approaches. Section 7 concludes with further discussion of our methodological and applied contributions.
2 Methods
2.1 Simplified causal framework
In this section, we introduce a simplified version of our causal identification strategy. While facilitating exposition, this simplicity comes at the expense of strong identifying assumptions, which we relax in Section 3. Let $Y_{it}$ be an outcome for observation $i$ at time $t$, with $i=1,\dots,n$ and $t=0,\dots,T$ and let $\bar{Z}_{it}$ denote the history of some random variable $Z$ from time $t=0,...,t$. We also let $A_{it}$ denote an indicator of a policy of interest being active for observation $i$ at time $t$, and define $Y_{it}(\bar{a})$ as the potential outcome that would be observed if the treatment history $\bar{A}_{it}$ were fixed to level $\bar{a}$. Our goal is to estimate the average treatment effect at time $t$, given by $\Delta_{t}=\mathbb{E}[Y_{it}(1)-Y_{it}(0)]$, where for now we only focus on treatment at time $t$. The challenge is that we do not observe $A_{it}=1$ and $A_{it}=0$ for all individuals - in fact, we never observe both. We therefore must make assumptions tying these counterfactual quantities to the observed data distribution. We consider more complex estimands that average over multiple policy years or depend on prior treatment values in Section 3. This choice of estimand, an average treatment effect, differs from the typical estimand identified by synthetic controls and difference-in-difference estimators, which estimate a treatment effect on the treated. We further discuss differences between these approaches further in Section 4.1.
To identify this effect, we first make the following consistency assumption,
Assumption 1 (Consistency).
$\bar{A}_{it}=\bar{a}\implies Y_{it}=Y_{it}(\bar{a}).$
Consistency states that the observed outcome under the observed policy history $\bar{A}_{it}=\bar{a}$ is equal to the potential outcome under that same policy history. This would be violated if there were multiple versions of the policy or when observed outcomes depend on other units’ policy statuses, sometimes referred to as “interference” or “spillover.” We next assume no carryover effects:
Assumption 2 (No carryover effects).
For any treatment history $\bar{a}$, $Y_{it}(\bar{a})=Y_{it}(a_{t})$
No carryover effects limits the dependence of the potential outcomes on the treatment history to only the current treatment status. In other words, lagged treatments have no impact on the current outcomes. Nevertheless, this assumption places no restrictions on how treatment effects may vary over time. This assumption is common in the policy evaluation literature when using methods such as difference-in-differences and synthetic controls.
We next assume ignorability, which states that treatment assignment $A_{it}$ is effectively randomized with respect to the potential outcomes in time period $t$ given $Y_{i,t-1}(0)$, the previous potential outcome absent treatment, which we will call the prior control counterfactual.
Assumption 3 (Ignorability).
$Y_{it}(a)\perp A_{it}\mid Y_{i,t-1}(0),\quad a=0,1.$
In contrast to standard ignorability assumptions, we assume that randomization occurs with respect to a potentially unobserved quantity – the potential outcome under control in time $t-1$, which is unobserved for those who had already been treated. We provide a more complete discussion of this identifying assumption in comparison to other assumptions in Sections 4.1 and 4.2. We also relax this assumption to control for additional prior counterfactuals and/or covariates in Section 3. To facilitate controlling for this prior counterfactual, we next assume treatment effects are constant across units:
Assumption 4 (Constant treatment effects).
$Y_{it}(a)=Y_{it}(0)+a\Delta_{t},\quad a=0,1.$
Constant treatment effects implies that the difference between the two counterfactuals is a fixed quantity at $t$, and it precludes any kind of random heterogeneity or heterogeneity that depends on unobservable quantities. This assumption may seem quite strong, particularly if certain units are expected to respond to treatment differently. We therefore explore violations of this assumption and how they may impact identification, along with relaxations of this assumption, in Section 2.3. Finally, we assume positivity.
Assumption 5 (Positivity).
For some $\epsilon>0$ and $e(y)=P\left\{A_{it}=1\mid Y_{i,t-1}(0)=y\right\}$,
$P\left[\epsilon<e\{Y_{i,t-1}(0)\}<1-\epsilon\right]=1.$
This requires that we have positive probability of observing either policy status conditional on the prior counterfactual under control. These core identifying assumptions allow us to write down our estimand as a function of the observed data distribution.
Theorem 1.
Under Assumptions 1-5,
$\Delta_{t}$ $=\mathbb{E}[\mathbb{E}[Y_{it}\mid A_{it}=1,Y_{i,t-1}(0)]-\mathbb{E}[Y_{it}\mid A_{it}=0,Y_{i,t-1}(0)]]$
$=\mathbb{E}\left[\mu_{1t}(Y_{i,t-1}-A_{i,t-1}\Delta_{t-1})-\mu_{0t}(Y_{i,t-1}-A_{i,t-1}\Delta_{t-1})\right],$ (1)
where $\mu_{at}(y)=E(Y_{it}\mid A_{it}=a,Y_{i,t-1}(0)=y)$. Thus, identification of the treatment effect at time $t$ requires identification of the treatment effect at time $t-1$.
2.2 Estimation
Here we describe approaches to estimating $\Delta_{t}$. We make one final assumption that facilitates the specification of estimators for the average treatment effect.
Assumption 6 (Time homogeneity).
For some $\Delta$, we have $\Delta_{t}=\Delta,\forall t.$
The time homogeneity assumption ensures that $\Delta_{t-1}$ (which we require for identification of $\Delta_{t}$) is identical to $\Delta_{t}$. We now review a few approaches to estimating $\Delta$. Under Assumptions 1-6, we can proceed by solving the following equation
$\psi_{1}=n^{-1}\sum_{i=1}^{n}Y_{it}\left[\frac{A_{it}}{e(Y_{i,t-1}-A_{i,t-1}\Delta)}-\frac{1-A_{it}}{1-e(Y_{i,t-1}-A_{i,t-1}\Delta)}\right]-\Delta=0.$
We cannot directly solve this without specification of $e(\cdot)$, so one could specify a parametric model $g\left\{e(Y_{i,t-1};\Delta)\right\}=\beta_{0}+\beta_{1}\left(Y_{i,t-1}-A_{i,t-1}\Delta\right)$ for a known link function $g(\cdot)$, and these two sets of parameters can be solved jointly using M-estimation theory.
Alternatively, we can specify a model for the conditional expectation function in equation (1), such as a debiased linear autoregressive model (Benjamin, Rigby and Stasinopoulos, 2003; Zeger and Qaqish, 1988) (L-DAM):
$Y_{it}$ $=\alpha+\delta(Y_{i,t-1}-\Delta A_{i,t-1})+\Delta A_{it}+\epsilon_{it}$ (2)
where we assume that the errors $\epsilon_{it}$ are conditionally mean zero with zero correlation over time. The L-DAM is essentially a simple linear parameterization of the causal assumptions presented above. We may once again use maximum likelihood to obtain estimates and inference for this model using M-estimation. Alternatively, we may specify priors for the parameters in (2) and use Bayesian inference, the approach we adopt in the rest of the paper.
These approaches could easily be adapted if Assumption 6 is implausible, or if time-specific effects are of particular interest. However, Assumption 6 connects our estimation approach to the traditional L-DAM in (2). Also, in many studies of state policies, at any given time, because data are limited, making assumptions like Assumption 6 that reduce the number of parameters to estimate may strike a favorable bias-variance trade-off.
2.3 Impacts of treatment effect heterogeneity
Assumption 4, which enforces constant treatment effects across units, is an especially strong assumption required for our estimation strategy. Under our modeling assumptions, it requires that $Y_{i,t-1}-\Delta A_{i,t-1}$ is equal to $Y_{i,t-1}(0)$. In the presence of treatment effect heterogeneity, this equivalence will not hold, leading standard estimation strategies to have both biased effect and variance estimates. Moreover, as heterogeneity increases, the degree of this bias increases. Note that there are many ways that ignoring treatment effect heterogeneity can negatively impact results that are not unique to our setting. However, because this particular form of misspecification – where we use estimated proxies to control for confounders – is unique to our identification strategy, we focus our discussion in this section on the impact of this misspecification, particularly with respect to the bias of the effect estimates.
2.3.1 Treatment effect heterogeneity as measurement error
Suppose Assumption 4 does not hold. Define $\Delta_{it}=Y_{it}(1)-Y_{it}(0)$ as the unit-level treatment effect for unit $i$ at time $t$ such that $\Delta=\mathbb{E}(\Delta_{it})$. We can re-write (2) to obtain
$Y_{it}$ $=\alpha+\delta(Y_{i,t-1}(0)+A_{i,t-1}\Delta_{i,t-1}-A_{i,t-1}\Delta)+\Delta A_{it}+\epsilon_{it}$
$=\alpha+\delta(Y_{i,t-1}(0)+U_{i,t-1})+\Delta A_{it}+\epsilon_{it}.$ (3)
In other words, adjusting for $Y_{i,t-1}-A_{i,t-1}\Delta$, our proxy of $Y_{i,t-1}(0)$, can be equivalently thought of as adjusting for a noisy version of $Y_{i,t-1}(0)$ subject to measurement error. An important parameter therefore becomes $\text{Var}(U_{i,t-1})$, which controls the magnitude of the heterogeneity and measurement error. Clearly, if $\text{Var}(U_{i,t-1})=0$ and we have no heterogeneity, then Assumption 4 holds, and Theorem 1 states that we can identify $\Delta$.
Throughout the rest of this section, we focus on a simplified $\widehat{\Delta}$ that comes from estimating the model in (3), ignoring the fact that $U_{i,t-1}$ is a function of $\Delta$, and focusing on the impact of this sort of measurement error in this setting. [^1] We find identical results when fitting the model in (2), so we proceed with the simpler setting that has easier to understand mathematical derivations. Measurement error in a confounder has been studied previously (Carroll, 1998), and it was shown that if $U_{i,t-1}$ is mutually independent of $Y_{i,t-1}(0),A_{it}$, and $Y_{it}$, then
$\mathbb{E}(\widehat{\Delta})=\Delta+\delta\gamma_{A}\bigg{(}\frac{\text{Var}(U_{i,t-1})}{\text{Var}(U_{i,t-1})+\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}}\bigg{)},$ (4)
where $\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}$ is the residual variance from regressing $Y_{i,t-1}(0)$ against $A_{it}$, and $\gamma_{A}$ is the coefficient for $A_{it}$ in that regression. If we did not adjust for $Y_{i,t-1}(0)$ at all, we would have $\mathbb{E}(\widehat{\Delta})=\Delta+\delta\gamma_{A}$. This shows that adjusting for an error prone version of $Y_{i,t-1}(0)$ (instead of $Y_{i,t-1}(0)$ itself) biases estimates towards what we would obtain without adjusting for the confounder at all, and the degree of this bias depends on the relative size of $\text{Var}(U_{i,t-1})$ and $\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}$. [^2] In state-level studies of opioid policies, we expect the treatment effect $\Delta$ to be relatively small in magnitude relative to the overall degree of variation in the outcomes, leading this bias to be small. Proposition 1 formalizes reasoning that may be used to justify the proposed estimator.
Proposition 1.
Under Assumptions 1, 3, 5, 6 and assuming model (3) holds, if $U_{i,t-1}$ is mutually independent of $Y_{i,t-1}(0),A_{it}$, and $Y_{it}$, and there exists a $k>0$ such that $\frac{\text{Var}(U_{i,t-1})}{\sigma_{Y_{i,t-1}(0)\mid A_{it}}}<k$, then we have
$\left|\mathbb{E}(\widehat{\Delta})-\Delta\right|\leq|\delta\gamma_{A}|\frac{k}{k+1}$
Proposition 1 shows that only large amounts of heterogeneity would lead to substantial bias in treatment effect estimates when the measurement error follows a classical structure and is independent of the treatment, outcome, and potential outcomes. However, this may not hold in many settings. We therefore next discuss the impacts of differential measurement error.
2.3.2 Differential measurement error and sources of heterogeneity
If the measurement error terms $U_{i,t-1}$ are correlated with any of $A_{it}$, $Y_{it}$, or $Y_{it}(0)$ then the bias expression in (4) no longer holds. The exact bias expression in this setting can be found in Appendix B, where we discuss the impacts of differential measurement error from treatment effect heterogeneity in greater detail. While we leave most details to the Appendix, we focus here on a particular form of differential measurement error that is most likely to occur and is most problematic.
The most likely reason in our context for differential measurement error arises due to correlation between $U_{i,t-1}$ and $Y_{it}$. When $U_{i,t-1}$ is high, unit $i$ was treated and had a large treatment effect at time $t-1$. When unit-level treatment effects are smooth over time, a large value of $U_{i,t-1}$ likely correlates with a high value of $Y_{it}$ due to the large unit-level treatment effect. Applying the formula from Appendix B to the situation where $U_{i,t-1}$ is correlated with $Y_{it}$, but not $A_{it}$, the estimate of the treatment effect has expectation
$\mathbb{E}(\widehat{\Delta})$ $=\frac{C_{1}+C_{2}+C_{3}}{\text{Var}(A_{it})-\frac{\text{Cov}^{2}(A_{it},Y_{i,t-1}(0))}{\text{Var}(Y_{i,t-1}(0)+U_{i,t-1})}}$
$=\Delta+\delta\gamma_{A}\bigg{(}\frac{\text{Var}(U_{i,t-1})}{\text{Var}(U_{i,t-1})+\sigma_{Y_{i,t-1}(0)\mid A_{it}}^{2}}\bigg{)}+\frac{C_{3}}{\text{Var}(A_{it})-\frac{\text{Cov}^{2}(A_{it},Y_{i,t-1}(0))}{\text{Var}(Y_{i,t-1}(0)+U_{i,t-1})}},$
where $C_{1}=\text{Cov}(A_{it},Y_{it})$,
$C_{2}=-\frac{\text{Cov}(A_{it},Y_{it}(0))\text{Cov}(Y_{it},Y_{it}(0))}{\text{Var}(Y_{i,t-1}(0)+U_{i,t-1})},\quad C_{3}=-\frac{\text{Cov}(A_{it},Y_{it}(0))\text{Cov}(Y_{it},U_{i,t-1})}{\text{Var}(Y_{i,t-1}(0)+U_{i,t-1})}.$
Notice that when the measurement error is classical and $\text{Cov}(Y_{it},U_{i,t-1})=0$, then $C_{3}=0$ and we revert back to the value in (4). Now we can examine whether $C_{3}$ is small (or large) in magnitude compared with $C_{1}$ and $C_{2}$ to examine whether the fact that the measurement error is differential plays a significant role in the magnitude of bias. We can see that the ratio of $C_{3}$ and $C_{2}$ can be written as,
$\frac{\text{Cov}(Y_{it},U_{i,t-1})}{\text{Cov}(Y_{it},Y_{it}(0))}=\frac{\text{Corr}(Y_{it},U_{i,t-1})\sqrt{\text{Var}(U_{i,t-1})}}{\text{Corr}(Y_{it},Y_{it}(0))\sqrt{\text{Var}(Y_{it}(0))}}.$
We typically expect $\text{Corr}(Y_{it},Y_{it}(0))$ to be high and likely larger in magnitude than $\text{Corr}(Y_{it},U_{i,t-1})$. Therefore, when $\text{Var}(U_{i,t-1})$ is much less than $\text{Var}(Y_{it}(0))$, differential measurement error will contribute little to the overall bias, and we will obtain bias similar to the expression given in Proposition 1. Moreover, this bias is expected to be small absent large effect heterogeneity.
2.3.3 Addressing treatment effect heterogeneity
We propose two strategies to proceed when estimating treatment effects using the autoregressive model (equation 2) in the presence of effect heterogeneity: in short, we can either ignore or model the heterogeneity. In most instances, it will be reasonable to assume that the hetereogeneity, and therefore the induced bias, is small. We explore this further in Section 5, where we find that heterogeneity of the treatment effect only impacts inference on the average treatment effect when the degree of heterogeneity is very large.
For cases where we cannot assume that effect heterogeneity is sufficiently small, we can modify the autoregressive framework to model the effect heterogeneity. For example, if $\Delta_{it}=h(X_{i},t)$, where $X_{i}$ are observed covariates, then we can easily specify a parametric form for $h(\cdot)$. We can then adjust for $Y_{i,t-1}-A_{i,t-1}h(X_{i},t-1)$, which should be a good proxy for the true $Y_{i,t-1}(0)$ for a reasonably well-specified $h(\cdot)$ function, alleviating effect heterogeneity induced biases. However, in some applications, we may have little data to estimate treatment effects, and this additional modeling may create too many parameters and lead to overly variable effect estimates. By contrast, in settings where we have relatively large sample sizes, this may be a perfectly reasonable strategy. On the other hand, if the heterogeneity is not a function of $X_{i}$, then we cannot use this approach and may instead rely on the results of the previous sections that argue this bias is likely small in magnitude. For example, if the heterogeneity takes the form $\Delta_{it}=h(V_{i})$ for missing variables $V_{i}$, then the resulting measurement error will likely be correlated with $Y_{it}$, as previously discussed.
One aspect unique to panel data that is crucial for this discussion is that we can partially assess the degree of heterogeneity empirically, and therefore whether the assumption that it does not significantly bias our results is reasonable. For treated unit-time pairs, we observe $Y_{it}=Y_{it}(0)+\Delta_{it}$. Assuming that the individual treatment effects $\Delta_{it}$ are not strongly and negatively correlated with $Y_{it}(0)$, then we would expect $\text{Var}(Y_{it}(0))\leq\text{Var}(Y_{it})$ for treated units at time $t$. We can estimate the variance of $Y_{it}$ during treated time periods and compare that with the variance of $Y_{it}$ during control periods, possibly adjusting for time-varying trends in variability. Similar ideas have been applied in randomized trials to test for heterogeneity (Mills et al., 2021). The idea is that a larger estimated outcome variance in the treated time periods would suggest heterogeneous treatment effects and would therefore potentially be a larger concern. This is not a formal test of whether heterogeneity is present, because these variances can be the same even in the presence of heterogeneity when $Y_{it}(0)$ and $\Delta_{it}$ are negatively correlated; nevertheless, this provides some additional information about the plausibility of the constant treatment effects assumption.
3 Relaxing assumptions and extensions
Up to this point we have focused on a simple set of identification assumptions to provide a straightforward and intuitive explanation of our identification and estimation strategies. While useful for exposition, these assumptions were quite strong, and can be weakened substantially. In this section, we relax these assumptions and show how these relaxations affect our proposed estimation strategy.
3.1 Additional time dependence
The main extensions we present allow for greater dependence of the variables over time, which allows us to relax the no carryover effects and ignorability assumptions. First consider no carryover effects: in many policy settings, treatment in earlier time periods likely affects outcomes at time $t$, and therefore, the effects of entire treatment sequences are of interest. To incorporate these effects we first redefine our causal estimand, now given by $\Delta_{t}(\bar{a})=\mathbb{E}[Y_{it}(\bar{a})-Y_{it}(\bar{0})].$ This parameter represents the effect of a particular treatment sequence compared against no treatment at time $t$. While in principle this allows for effects of all prior treatments, in practice, we make the assumption that this is limited to only the most recent $l$ time points.
Assumption 7 ($l$-limited carryover effects).
Let $\bar{a}$ and $\bar{a}^{\prime}$ be two treatment sequences such that $a_{t}=a_{t}^{\prime}$, $a_{t-1}=a_{t-1}^{\prime},\dots,a_{t-l}=a_{t-l}^{\prime}$, then we have $Y_{it}(\bar{a})=Y_{it}(\bar{a}^{\prime}),\ \forall\bar{a}.$
Assumption 7 states that the potential outcomes for an entire treatment sequence only depend on the previous $l$ treatment values. Notice that when $l=0$ we return to the no carryover effects assumption. For simplicity, when we refer to $\bar{a}$ or $\bar{A}_{it}$ in what follows, we only refer to the treatment sequence for the previous $l$ time periods.
Second, we relax Assumption 3 to condition on $k$ prior time period’s counterfactuals as opposed to only the single time period’s counterfactual outcome used previously. This relaxation increases the plausibility of the assumption, particularly in settings where we worry about the dependence of the outcomes over time. However, in order to relax this assumption, we must also strengthen the positivity assumption (Assumption 5) to include these additional time-periods. We can then incorporate these relaxations into our estimation strategy by fitting the following version of the L-DAM:
$Y_{it}$ $=\alpha+\sum_{b=1}^{k}\left\{\delta_{b}\left(Y_{i,t-b}-\sum_{z=0}^{l}\theta_{z}A_{i,t-b-z}\right)\right\}+\sum_{z=0}^{l}\theta_{z}A_{i,t-z}+\epsilon_{it}$
It is easy to see how this model incorporates these extensions. First, lagged policy effects are now given by the $\theta_{z}$ parameters. Additionally, the model includes linear adjustments for estimates of $Y_{i,t-1}(0),\dots,Y_{i,t-k}(0)$ instead of estimates of just $Y_{i,t-1}(0)$.
3.2 Time-varying covariates
We can also relax our ignorability assumption to allow for both time-invariant and time-varying covariates. Assume, for example, that covariates $X_{it}$ are determined prior to $Y_{it}$. Then we could assume that ignorability holds conditional on covariates and $k$ lagged potential outcomes absent treatment $(X_{it}$, $Y_{it-k}(0),...,Y_{it-1}(0))$. However, we must again change the positivity requirement to further condition on these quantities. One can then incorporate these same covariates into the L-DAM as:
$Y_{it}$ $=\alpha+\sum_{b=1}^{k}\left\{\delta_{b}\left(Y_{i,t-b}-\sum_{z=0}^{l}\theta_{z}A_{i,t-b-z}\right)\right\}+\sum_{z=0}^{l}\theta_{z}A_{i,t-z}+\Gamma^{\top}X_{it}+\epsilon_{it}$
When the covariates $X_{it}$ are time-varying, implicit in this specification are that the covariates are exogenous, or unaffected by treatment. That is, we further assume,
Assumption 8 (Covariate exogeneity).
$X_{it}(\bar{a})=X_{it}(\bar{a}^{\prime})=X_{it},\qquad\forall\bar{a},\bar{a}^{\prime}.$
This assumption is frequently made in the applied policy literature, including when using two-way fixed effects models with time-varying covariates (Wooldridge, 2023). In cases where this assumption is implausible, one can instead simply adjust for pre-treatment lags of the covariates, which, absent anticipatory effects, are unaffected by treatment by design.
3.3 Multiple policies
A separate but important type of time-varying variable we may also observe are other policy indicators $M_{it}$. For example, state legislatures may consider multiple policies to tackle opioid-related outcomes, including both naloxone access laws and Medicaid prescription coverage policies. In this situation, we may wish to know the effect of one policy, holding the other policy fixed at some value. In the language of potential outcomes, we are interested in a so-called “controlled direct effect” (Nguyen, Schmid and Stuart, 2021).To be precise, assume that in each time period states may assign either policies $A$ or $M$, but we are primarily interested in the effects of policy $A$. One relevant causal quantity would then be $\psi_{t}(\bar{a},\bar{m})=\mathbb{E}[Y_{it}(\bar{a},\bar{m})-Y_{it}(\bar{0},\bar{m})]$, where potential outcomes are now indexed by both treatment values. Identifying this quantity in terms of the observed data distribution requires analogues to Assumptions (1)-(5). We formally describe these assumptions in Appendix A; however, intuitively we can view $(A,M)$ as a multi-valued treatment randomly assigned conditional on the prior potential outcomes absent any policy. [^3] We can allow that the randomization for $M$ is conditional $A$ or vice versa. In much of the mediation literature it is assumed that $A$ is determined prior to $M$. We instead make no assumption on their causal ordering, at the price of a slightly stronger conditional independence assumption. Under these assumptions we can identify the causal estimand in terms of the observed data distribution as an average of a conditional expectation function, described in Appendix A. For continuous outcomes we can again parameterize a model for this function using an L-DAM:
$Y_{it}$ $=\alpha+\sum_{b=1}^{k}\left\{\delta_{b}\left(Y_{i,t-b}-\sum_{z=0}^{l}[\theta_{z}A_{i,t-b-z}+\zeta_{z}M_{i,t-b-z}]\right)\right\}+\sum_{z=0}^{l}[\theta_{z}A_{i,t-z}+\zeta_{z}M_{i,t-z}]+\epsilon_{it}$
Under this model, $\theta_{z}$ represents the direct effect of policy indicator $A_{t-z}$ on the outcome $Y_{it}$ holding the policy sequence $\bar{M}_{it}$ constant. As before, we can expand this model to include covariates and more complicated forms of treatment effects that allow, for example, effects due to interactions between policies $A$ and $M$. However, in small data settings, it may not be feasible to estimate these interactions with adequate precision.
3.4 Nonlinear extensions
Linear models may be inappropriate for modeling the conditional expectation function for certain outcomes. For example, in opioid studies, we are often faced with count data, where a log-link function may be more appropriate. We can continue to invoke Assumptions (1)-(5) above, but in place of assumption (4), we may instead assume constant multiplicative treatment effects:
Assumption 9 (Constant multiplicative treatment effects).
Letting $\eta_{t}(\bar{a};\theta_{t})$ be a constant treatment effect for treatment sequence $\bar{a}$ that is a function of parameters $\theta_{t}$ such that $\eta_{t}(\bar{0};\theta_{t})=0$, we have $\frac{Y_{it}(\bar{a})}{Y_{it}(\bar{0})}=\exp(\eta_{t}(\bar{a};\theta_{t})).$
Analogous to the L-DAM, we can then parameterize the conditional expectation function with respect to a log-link and assuming $\eta_{t}(\bar{a};\theta_{t})=\theta^{\top}\bar{a}$:
$\log(\mathbb{E}[Y_{it}\mid\bar{Y}_{i,t-1},\bar{A}_{it}])$ $=\alpha+\sum_{b=1}^{k}\left\{\delta_{b}\left(\log(Y_{i,t-b})-\sum_{z=0}^{l}\theta_{z}A_{i,t-1-z}\right)\right\}+\sum_{z=0}^{l}\theta_{z}A_{i,t-z}.$
Under constant multiplicative treatment effects, we have that
$Y_{it}=Y_{it}(\bar{0})\exp(\eta_{t}(\bar{A}_{it};\theta_{t}))=Y_{it}(\bar{0})\exp(\theta^{\top}\bar{A}_{it}),$
which implies that $\log(Y_{it}(\bar{0}))=\log(Y_{it})-\theta^{\top}\bar{A}_{it}$. This reveals that in the autoregressive model, we are controlling for the log of counterfactual outcomes, given by $Y_{i,t-b}(\bar{0})$. Additionally, the estimand of interest may change to one on the multiplicative scale, depending on the application of interest, leading to
$\tilde{\Delta}_{t}(\bar{a})=\frac{\mathbb{E}[Y_{it}(\bar{a})]}{\mathbb{E}[Y_{it}(\bar{0})]}.$
The log-linear model and multiplicative estimand add no additional complications. We may continue to use our model to obtain estimates of $\mathbb{E}[Y_{it}(\bar{a})]$ and $\mathbb{E}[Y_{it}(\bar{0})]$ by marginalizing the model over the empirical covariate distribution, setting $\bar{A}=\bar{a},\bar{0}$, and using the averages of these predictions to compute the estimates of causal contrasts. One potential issue arises when there are zero counts in the data (as $\log(0)$ is undefined). In these settings, we can adjust for a different function of the lagged outcomes other than the log, possibly using basis function expansions to let the data inform the shape of this function.
4 Connections to existing identification strategies
In this section we briefly compare our identification strategy to other strategies commonly used in causal inference with panel data. We first compare our approach to difference-in-differences and synthetic controls, commonly used in applied policy applications, and we then compare it to the sequential g-formula, which identifies causal quantities under sequentially randomized trials and is a more commonly used approach in epidemiologic studies.
4.1 Difference-in-differences and synthetic controls
As mentioned before, difference-in-differences and synthetic controls conventionally target the treatment effect on the treated (ATT), while our identification strategy can be used to target the ATT or the average treatment effect (ATE). Because the potential outcomes under treatment are fully observed, for DiD and synthetic controls analyses, the only quantity requiring causal identification is the mean potential outcomes absent treatment among the treated group. Synthetic controls do this by assuming ignorability of the potential outcomes absent treatment conditional on pre-treatment outcomes. [^4] Alternatively, synthetic controls have frequently been motivated by assuming that the potential outcomes absent treatment follow a linear factor model (Abadie, Diamond and Hainmueller, 2010). By contrast, DiD assumes “parallel-trends:” that the mean trends in the untreated group equal the counterfactual trends absent treatment in the treated group. This assumption allows for selection into treatment to be based on an unmeasured variable causing a time-invariant difference between the mean outcomes in the treatment and control groups.
To compare these assumptions to ours, consider the canonical DiD setting with two time-periods ($t=1,2$), where treatment only occurs in the second time-period. In this case, our ignorability assumption simply states that treatment is randomized with respect to the potential outcomes in the second time-period conditional on the first-period outcomes. Moreover, this assumption implies the ignorability assumption justifying synthetic controls Abadie, Diamond and Hainmueller (2010); Ben-Michael, Feller and Rothstein (2021). One key difference, is that our assumption is stronger: we require ignorability for both potential outcomes – $Y_{i2}(1),Y_{i2}(0)$ – while synthetic controls only requires an ignorability assumption with respect to $Y_{i2}(0)$. Beyond this, any comparison between conditional randomization and parallel trends in this setting immediately applies to our approach. In a more general setting, our approach essentially extends the synthetic controls ignorability assumption by imputing the counterfactuals absent treatment and conditioning on these quantities to obtain causal identification. In other words, while synthetic controls only conditions on the potential outcomes absent treatment prior to the implementation of treatment, our approach allows the analyst to condition on any subset of potential outcomes absent treatment up to the previous time-period, regardless of whether treatment occurred in the previous time-period. As a result, our ignorability assumption may be seen as an extension of an ignorability assumption underlying synthetic controls.
One key benefit of our approach relative to DiD or synthetic controls is the ability to estimate causal effects in the presence of dynamic policy landscapes where multiple policies change over time. In DiD or synthetic controls applications, the presence of other policy changes may be problematic. In DiD contexts, additional policy changes are essentially time-varying covariates. Because these policy changes likely differ in distribution between the treatment and control groups, these changes threaten the parallel-trends assumption and potentially confound DiD estimates (Zeldow and Hatfield, 2021). While identification may be recovered by conditioning on these policy changes as time-varying covariates, this requires a further assumption that the primary policy change of interest has no effect on other policy changes, which may not hold in practice. For synthetic controls applications, the presence of multiple changing policies does not necessarily threaten causal identification. However, the effect estimates will capture both a direct effect of the relevant policy on the outcome, and an indirect effect via the induced policy changes. As a result, the effect estimates capture effects of multiple policies. When the indirect effects are thought to be small relative to the direct effects, this may not be a problem; otherwise, these changes make the estimates harder to interpret. By contrast, our approach, when applying the extension to handle multiple policies, enables the identification of the direct effects of the policies, partialling out the effects of other policy changes. Moreover, unlike the DiD context, we require no assumption about the dependence of the policies to obtain identification.
In summary, our approach relies on an ignorability assumption that is similar to an ignorability assumption that may motivate synthetic controls; in fact, these assumptions may be identical in settings with only one treated period. The key difference is that our approach allows conditioning on post-treatment potential outcomes absent treatment, whereas synthetic controls only conditions on pre-treatment outcomes. Finally, in the presence of multiple policy changes, our approach can identify the direct effects a single policy change, whereas identification itself is threatened in a DiD context, and synthetic controls estimate will capture both direct and indirect effects, which may not be desirable.
4.2 The sequential g-formula
We now compare the identification result in Theorem 1 with the sequential g-formula (Gill and Robins, 2001). A more common approach in the epidemiology literature, the sequential g-formula relies on the so-called “sequential ignorability” assumption, which essentially states that in every time period, treatment is randomly assigned conditional on all prior outcomes, treatments, and covariates.  [^5] Sequential ignorability states that for all time-periods $t$, $Y_{it}(\bar{a}_{t})\perp A_{it}\mid\bar{Y}_{i,t-1},\bar{X}_{i,t-1},\bar{A}_{i,t-1}$., In combination with the corresponding positivity assumptions, sequential ignorability applied to our setting yields the identifying expression (the sequential g-formula),
$\mathbb{E}[Y_{it}(\bar{a})]$ $=\int_{Y_{0}}...\int_{Y_{T-1}}\mathbb{E}[Y_{T}\mid\bar{Y}_{T-1},\bar{A}_{T-1}=\bar{a}_{T-1}]\prod_{t=0}^{T-1}dP(Y_{t}\mid\bar{Y}_{t-1},\bar{A}_{t-1}=\bar{a}_{t-1}).$ (5)
This expression is far more complex than the functional given in Theorem 1, which expresses our identified causal parameter in terms of a single averaged conditional expectation function. [^6] In practice, to estimate equation (5) structure is often placed on $\mathbb{E}[Y(\bar{a}_{t})]$; such models are known as marginal structural models (Robins, Hernan and Brumback, 2000). We obtain a simpler expression because our ignorability assumption conditions on the counterfactual outcomes absent treatment rather than the observed outcomes. To see this, consider a simple case with three time-periods, illustrated in Figure 1 below.
[…]
Figure 1: Conventional longitudinal causal structure
Ideally, we would hope to identify $\mathbb{E}[Y_{2}(\bar{a})]$ in terms of an averaged conditional expectation function. One possibility is the averaged autoregressive model,
$\mathbb{E}[\mathbb{E}[Y_{2}\mid Y_{1},Y_{0},A_{1}=a_{1},A_{2}=a_{2}]].$
However, equivalence between this expression and $\mathbb{E}[Y_{2}(\bar{a})]$ would require that $(A_{1},A_{2})\perp Y_{2}(a_{1}a_{2})\mid(Y_{0},Y_{1})$. Unfortunately, this will not hold under the causal structure depicted in Figure 1. The problem, as mentioned previously, is that $Y_{1}$ is a post-treatment outcome that acts as a collider between $A_{1}$ and $Y_{2}(a_{1}a_{2})$, since $Y_{1}$ lies on the causal pathway from $A_{1}$ to $Y_{2}$. Conditioning on $Y_{1}$ introduces a dependence between $Y_{2}(a_{1}a_{2})$ and $A_{1}$ that otherwise may not exist, violating the required ignorability assumption noted above.
By contrast, our ignorability assumption conditions on $(Y_{1}(00),Y_{0}(00))$, variables that do not lie on the causal pathway from $A_{1}$ to $Y_{2}$. Figure 2 depicts a causal diagram illustrating the assumed relationships between variables, where blue arrows indicate deterministic relationships (notice that the potential outcomes $Y_{t}(a_{1}a_{2})$, omitted from this diagram, are a deterministic function of $Y_{t}(00)$ by assumption). Figure 2 implies the conditional independence statement $(A_{1},A_{2})\perp Y_{2}(a_{1}a_{2})\mid Y_{1}(00),Y_{0}(00)$. By conditioning on the unobserved potential outcomes absent treatment rather than the observed outcomes, our approach does not introduce collider bias, and we are therefore able to achieve causal identification of an averaged conditional expectation function.
[…]
Figure 2: Assumed longitudinal causal structure
However, this identification strategy comes with a cost: because we do not necessarily observe these counterfactual quantities required for ignorability, we instead rely on structural assumptions that parameterize the treatment effects as a function that must be estimated. By contrast, equation (5) makes no such restriction. As a result, our approach allows for simpler estimation methods, but at the cost of stronger modeling assumptions.
5 Simulation studies
We evaluate the performance of the proposed approach against state-of-the-art methods for panel data causal inference in a simulation study. Rather than simulating outcomes from a data generating process we choose, we instead utilize existing outcome data from our application to ensure that the results are more relevant to our setting. Specifically, we let our observed outcome data represent $Y_{it}(0)$, and then we simulate treatment effects to obtain $Y_{it}(1)$. Our data consist of 51 units and 16 time periods, and we choose 20 units to be treated in all simulations. Our observed outcomes are the number of opioids prescribed by each state (and Washington DC) and year. We do not vary the outcome values across simulations, but instead vary which units are treated, when they are treated, and the heterogeneity of the causal effect (when applicable). Specific details about the simulation scenarios can be found in Appendix D, though we summarize the key points here. We let treatment status at time $t$ depend on $Y_{i,t-1}(0)$, and follow a staggered adoption setting where a unit remains treated once they are treated. We vary the degree of confounding by $Y_{i,t-1}(0)$ into three categories we refer to as no confounding, moderate confounding, and high confounding. We set a small average treatment effect and vary the degree of heterogeneity of this effect across units, split into categories denoted by no heterogeneity, moderate heterogeneity, and high heterogeneity. The moderate and high heterogeneity settings represent scenarios where the probability that a unit has an individual treatment effect either twice as large, or the opposite sign, as the average treatment effect is set to be 0.1 and 0.33, respectively. It is also important to emphasize that the only assumption placed on simulated data that coincides with our modeling assumptions is that the effects are homogeneous with respect to time. Every other aspect of the simulated data comes from the real, observational data, and may or may not fit well into the autoregressive framework we presented.
5.1 Target estimand and competing estimators
We explore four different estimators in this manuscript: 1) The autoregressive model that assumes a linear autoregressive model on the rates of the outcome that incorporates two time period lags, 2) The two way fixed effects estimator, 3) the synthetic control estimator for staggered adoption developed in Ben-Michael, Feller and Rothstein (2022), and 4) the difference in difference estimator for staggered adoption developed by Callaway and Sant’Anna (2021). Note that we use the linear autoregressive model here that uses a normal approximation to the rates of the outcome instead of directly modeling counts as described in Section 3.4. In Appendix D we run the same simulations using negative binomial autoregressive models and find largely similar results. Further, to ensure a fair comparison across different estimation procedures, our target of inference throughout is the sample treatment effect on the treated:
$\frac{1}{n_{t}}\sum_{i\in\mathcal{N}_{t}}\sum_{m=0}^{2}\bigg{\{}Y_{i,T_{i}+m}(1)-Y_{i,T_{i}+m}(0)\bigg{\}}.$
This estimand targets the average treatment effect on the treated units in the first three time periods after policy adoption. For more details on how to perform inference on sample level estimands within our framework, see Appendix C or read Ding and Li (2018) for a general discussion. Most importantly, however, this estimand is easily estimated using any of the synthetic control, DiD, and autoregressive models we consider. Interpretation of the TWFE estimator may be complicated in this setting (Imai and Kim, 2021), but we include it due to its widespread use. We calculate standardized mean squared error (MSE), which in any simulation scenario is the MSE of an estimator divided by the minimum MSE across all estimators, leading to the optimal estimator having an MSE of 1. We also calculate 95% interval coverage rates and the power to detect a nonzero effect, which is calculated as the proportion of 95% intervals for which zero is not included and the interval correctly estimates the sign of the treatment effect.
5.2 Results on simulation using opioid prescribing data
Figures 3 - 6 display the results. The best performing estimators are the autoregressive model and synthetic control approach, followed by the difference in differences estimator, and then the TWFE estimator. The 95% interval coverage is good for the autoregressive, DiD, and SC approaches with the autoregressive approach only falling below 95% coverage with high heterogeneity or large amounts of confounding. The autoregressive approach has the best MSE in all scenarios with a standardized MSE of 1, and it is substantially more efficient in certain scenarios. This highlights the bias-variance trade-off inherent to using these models as they can be significantly more efficient than existing estimators that are not model based. This boost in efficiency is most notable in Figure 6, where we display power. Both the DiD and SC approaches have effectively no power to detect the nonzero treatment effects, while the autoregressive approach detects this effect in over half of the simulated data sets. This is crucial, because policy effects are typically very small in state-level opioid policies and it appears that existing approaches are not well-suited to detecting these effects. Another key takeaway from the simulation is that reasonable amounts of heterogeneity does not appear to negatively impact the performance of the autoregressive model, consistent with the results of Section 2.3 suggesting that only very large amounts of heterogeneity will lead to poor performance of these models.
Figure 3: Distribution of point estimates for simulations on opioid prescribing data.
Figure 4: 95% interval coverage for simulations on opioid prescribing data.
Figure 5: Mean squared error for simulations on opioid prescribing data. MSE values have been capped at 5 for easier comparison.
Figure 6: Power for simulations on opioid prescribing data.
6 Examining the impact of state policies on opioid prescriptions
We now examine the impact of four state-level policies on the number of opioid prescriptions per capita in each state over the years 2006-2021. We first provide background on the data sources used and the questions we set out to answer, and then utilize the proposed autoregressive model to estimate the effects of these policies.
6.1 Covariates, policies, and outcome of interest
To measure the key policies of interest, we utilized OPTIC-Vetted Policy Data Sets from the RAND-USC Opioid Policy Tools and Information Center (OPTIC) to provide us with high quality policy measures (RAND-USC Schaeffer Opioid Policy Tools and Information Center, 2024a, b, c). We choose to focus on four different policies that have been examined in prior research but are rarely considered jointly. Naloxone is a potential life saving treatment for opioid overdoses that works by reversing the effects of opioids (Cawley and Dragone, 2023; Chamberlain and Klein, 1994), though individuals face systemic stigma or lack access to traditional healthcare systems, making it difficult to obtain naloxone (Green et al., 2020; Smart and Grant, 2021). To address these barriers, every state has adopted some type of NAL by 2018 (Smart, Pardo and Davis, 2021). There are two broad classes of naloxone access laws (NALs Smart et al. (2023)): those which allow distribution through a standing or protocol order or those which allow distribution through a pharmacist prescriptive authority (Davis and Carr, 2015, 2017).
A third policy we study is prescription drug monitoring programs (PDMPs), which are electronic databases that track prescribed and dispensed controlled substances within a state (Schuler et al., 2020). By giving healthcare providers information about a patient’s prescription drug history, PDMPs may help identify patients who are at risk for opioid misuse or overdose, potentially increasing early intervention and support (Horwitz et al., 2018). Based on the literature on PDMPs, state laws that specify that prescribers must access the PDMP before writing a prescription to a new/existing patient tends to allow for more effective forms of PDMPs (Pacula and Stein, 2020; RAND-USC Schaeffer Opioid Policy Tools and Information Center, 2024b). Thus, we define our PDMP laws as those which have this must access component. Lastly, many states have legalized marijuana for medical purposes (Pacula et al., 2015). Medical marijuana laws have the potential to reduce opioid use and opioid-related harms by providing an alternative treatment option for chronic pain and other conditions that are commonly treated with opioids (Powell, Pacula and Jacobson, 2018). For all of these policies, we record the fraction of the year each state has such a law in place. For a summary of the prevalance of these laws, and how this varies over time, see Figure 7
Figure 7: Policy wheels showing the timeline of policy adoption
We examined the impact of these policies on annual CDC opioid prescribing rates per a population of 100,000 people from 2006-2021, constructed from the QuintilesIMS Transactional Data Warehouse (TDW). These data represent approximately 59,000 retail (non-hospital) pharmacies, which dispense nearly 88% of all retail prescriptions in the U.S. Opioid prescriptions were identified using the National Drug Code, while cough and cold formulations containing opioids and buprenorphine, an opioid partial agonist used for treatment of opioid use disorder as well as for pain, were not included. For rates, annual resident population denominator estimates were obtained from the Population Estimates Program, U.S. Census Bureau. An illustration of opioid prescription rates over time by state can be found in Appendix E. Finally, we included seven time-varying covariates in our models that include unemployment rates, median income, and poverty rates, as well as percentages of non-Hispanic white individuals, females, individuals under 20 years old, and individuals greater than 60 years old. These covariates were selected to capture a range of sociodemographic factors that might confound the effects of our policies on the outcomes.
6.2 Single policy analysis
First, we analyze the impact of each policy on prescribing separately, which facilitates comparison with existing approaches designed for one policy only. Additionally, our policy variables are proportions between 0 and 1, but for this analysis, we dichotomize policy variables based on whether the state has any law in place that year.
6.2.1 Model formulation and estimands of interest
Our estimand of interest is the sample average treatment effect on the treated used in Section 5, which represents the average impact of the policy on the treated units in the first 3 years it is enacted. We utilize four distinct estimators of this quantity: 1) a DiD estimator allowing for staggered adoption, 2) the augmented synthetic control approach that allows for staggered adoption, 3) the autoregressive model incorporating time-varying covariates, and 4) the autoregressive model without time-varying covariates. For the autoregressive models, we must specify the specific form of our model as well as the prior distributions to be used. We assume that $k=2$ and that a second order autoregressive model is sufficient, and we enforce that only the policy indicator at the most recent three time points affects the outcome. The autoregressive coefficients are assigned standard uniform prior distributions, while all regression parameters, including those that govern the treatment effect, are assigned noninformative normal priors centered at zero with a variance of 100, and the variance parameters are assigned inverse gamma priors with parameters 0.001 and 0.001.
6.2.2 Results
The results for each of the four policies under each of the four estimators is given in Table 1. Note there is no estimate for the augmented synthetic control estimator for the medical marijuana policy, as many states had already enacted this policy at the start of our study and therefore this policy is not amenable to synthetic controls. For both NAL policies as well as medical marijuana, there is a fairly consistent story across all estimators, which is that they do not have an impact on opioid prescriptions. All estimators estimate a negative effect of PDMP laws on prescriptions suggesting must access PDMP laws decrease opioid prescribing, with both the DiD and autoregressive models showing a statistically significant effect. In general across all four policies, the autoregressive models tend to have much smaller interval widths, highlighting the efficiency gains that come with this approach.
In addition to the sample average treatment effect that is averaged over the first three time periods, we also estimate a time specific sample average treatment effect that does not average over time periods. We show estimates of this quantity for the autoregressive model with covariates in Figure 8. We see no effect at any time period for NAL or medical marijuana laws, which is expected given the previous results. We see that the effect of PDMP laws becomes more pronounced after the first year. This could be due to the fact that we dichotomized the treatment variable, but in the first year in which it is enacted, it is commonly only enacted for a fraction of the year.
DiD Augmented SC AR AR no covariates
NAL protocol standing -0.50 (-3.06, 2.05) -0.20 (-6.03, 5.31) 0.04 (-0.92, 0.94) 0.02 (-0.89, 0.91)
NAL prescriptive authority 0.32 (-2.45, 3.09) 1.35 (-5.61, 8.11) -0.22 (-2.09, 1.6) -0.25 (-2.21, 1.59)
Must Access PDMP -4.77 (-8.87, -0.68) -3.86 (-9.77, 1.58) -2.27 (-3.7, -0.85) -2.29 (-3.72, -0.84)
Medical marijuana 1.34 (-1.23, 3.91) Not available -0.08 (-1.16, 0.99) -0.13 (-1.12, 0.87)
Table 1: Estimates and 95% intervals for the sample average treatment effect on the treated across all estimators.
Figure 8: Estimates, and corresponding 95% credible intervals, for the time-specific sample average treatment effect on the treated from the autoregressive model that includes covariates.
6.3 Examining multiple policies simultaneously
We can also use the proposed approach to estimate the effect of multiple policies simultaneously, allowing us to see whether certain policies are more (or less) effective if implemented in tandem with other policies. We no longer dichotomize treatment variables and instead use the proportion of that year that the policy was implemented as our main treatment variables. We use a similar model formulation as for the single policy case as we utilize $k=2$ autoregressive lags, though here we only allow the treatments at the current time point to affect outcome. This is done partially to reduce the complexity of the model in order to allow us to study multiple policies and their interactions simultaneously. For this analysis, we focus on two policies of interest: must access PDMPs and medical marijuana laws. Studying all four laws simultaneously is exceedingly difficult with only 51 units and leads to serious issues with positivity violations as there are combinations of the four policies for which we have no data. We focus on the must access PDMP and medical marijuana laws specifically, because they are the two most likely to affect opioid prescription rates. Naloxone access laws are primarily expected to reduce opioid mortality by intervening in cases of an overdose, rather than by reducing prescriptions.
The estimand of interest is similar to the one studied for single policy analyses, but now potential outcomes are denoted by $Y_{it}(a,m)$, which is the outcome we would observe for unit $i$ at time $t$ had they been exposed to levels of the two policies given by $a$ and $m$. We look at a sample average potential outcome given by $SAPO(a,m)=\frac{1}{n}\sum_{i=1}^{n}Y_{i,T_{i}}(a,m)$, for all combinations of $a$ and $m$. This shows the average outcome, across all states, for each combination of the two policies. Moving forward, the first policy will represent must access PDMP laws, while the second policy will represent medical marijuana laws.
6.3.1 Results
The estimates and corresponding 95% credible intervals for SAPO($a$, $m$) for the four combinations of policies is given in Figure 9. The values of SAPO(00, 00) and SAPO(00, $1$) are both high indicating that medical marijuana does little to reduce opioid prescriptions. We also see that SAPO($1$, 00) is much lower, which re-iterates what was seen in the single policy analysis, which is that must access PDMP laws alone are able to substantially reduce opioid prescriptions at the state level. Interestingly, the value of SAPO($1$, $1$) is moderately higher than SAPO($1$, 00), which indicates that the presence of medical marijuana policies can dampen the influence of must access PDMPs. To see this, we can also examine the posterior distribution of the coefficients corresponding to the policy variables in the autoregressive model. The treatment effect component of the model is given by $A\theta_{pdmp}+M\theta_{mm}+(A\times M)\theta_{int}$, and the posterior distribution of these three parameters can be seen in Appendix E. The interaction term has posterior density almost entirely greater than 0 (95% CI: 0.50, 4.89), highlighting that medical marijuana laws do temper the effect of the must access PDMP laws on opioid prescriptions.
Figure 9: Estimates of SAPO for opioid prescriptions under all combinations of PDMP and medical marijuana laws
7 Discussion
We examined the impact of four policies on opioid prescribing, highlighting meaningful evidence that must access PDMPs decreases opioid-prescribing rates within a state. We also found null effects showing that naloxone access laws and medical marijuana did not meaningfully reduce opioid prescriptions at the state level, and that medical marijuana may reduce the impact of must access PDMPs when enacted jointly. Taken together, our findings suggest that a combination of policies may be needed to effectively address the opioid crisis. In the end, the application must be interpreted with caution when considering some of the null findings for the other policies. The policies we examine were enacted to improve the lives of individuals in the United States, and even small effects of policies could translate to saving or helping thousands of individuals.
Our work also makes a strong methodological contribution. In particular, we propose a novel causal framework that provides formal justification for using debiased autoregressive models to estimate causal effects. While our causal assumptions make strong restrictions on effect heterogeneity, we show mathematically that the presence of effect heterogeneity likely only causes small bias on our effect estimates. Moreover, we show how our approach easily extends to allow for identification and estimation in the presence of multiple policy changes. By contrast, effect estimation in this context is more challenging for difference-in-differences and synthetic controls methods, as discussed above in Section 4.1. Finally, we provide additional simulation evidence showing that our proposed method outperforms existing methods in opioid-policy settings where we have small sample sizes with a changing policy landscape, consistent with previous studies (Griffin et al., 2021; Schell, Griffin and Morral, 2018; Griffin et al., 2023, 2022). More generally, the proposed debiased autoregressive models offer a promising method for policy researchers to evaluate the effects of social and health policies in multiple applied areas, including gun violence, opioids, and COVID-19 policy evaluations. By providing relatively low variance estimates in small data settings, this method allows for more precise causal inferences with relatively low bias.
On the other hand, this method is not without limitations. Our identification and estimation strategy requires making strong functional form assumptions on the treatment effect function itself, a problem that may be difficult to reason about in certain applications. Additionally, while we have shown that effect heterogeneity is likely only to induce small bias with respect to effect estimation in a highly simplified setting, we have not formally shown that these results extend in more complex settings where we condition on multiple pre-treatment outcomes, covariates, and attempt to evaluate the effects of multiple policies. Valuable future research may therefore consider how to conduct formal sensitivity analyses with respect to the required causal assumptions, and to formalize how heterogeneity may affect model-based estimates in more complex settings. Nevertheless, we believe that we have made strong progress in formally justifying the use of debiased autoregressive models for causal effect estimation and for further arguing for their utility in applied policy settings.
Acknowledgements
This research was financially supported through a National Institutes of Health (NIH) grant (P50DA046351) to RAND (PI: B.D.S.) as well as Arnold Ventures (AV). NIH and AV had no role in the design of the study, analysis, and interpretation of data nor in writing the manuscript. We wish to thank Andrew Morral, Megan Schuler, Bradley Stein, Rosalie Pacula, David Powell, and Evan Peet for comments on different iterations of this paper and/or the execution of the policy evaluation. We also thank the members of the OPTIC Advisory Board for their feedback on these analyses.
Appendix A Causal assumptions for multiple policies
We define our target parameter as the counterfactual quantity $\mathbb{E}[Y_{it}(\bar{a},\bar{m})]$. To identify this quantity in terms of the observed datam assume the following:
Assumption 10 (Consistency).
$\bar{A}_{it}=\bar{a},\bar{M}_{it}=\bar{m}\implies Y_{it}=Y_{it}(\bar{a},\bar{m}),\qquad\forall\bar{a},\bar{m}$
This assumption extends the previous assumption to account for multiple policies. Similar to before, this assumption precludes interference across states (or units) – that is, each states’ observed outcome is only a function of the policies implemented in that state.
Assumption 11 ($l$-limited carryover effects).
Let $\bar{a},\bar{m}$ and $\bar{a}^{\prime},\bar{m}^{\prime}$ be two treatment sequences such that $a_{t}=a_{t}^{\prime}$, $a_{t-1}=a_{t-1}^{\prime},\dots,a_{t-l}=a_{t-l}^{\prime}$, and similarly for $\bar{m},\bar{m}^{\prime}$. Then we have
$Y_{it}(\bar{a},\bar{m})=Y_{it}(\bar{a}^{\prime},\bar{m}^{\prime}),\qquad\forall\bar{a},\bar{m},\bar{a}^{\prime},\bar{m}^{\prime}.$
This again extends the previous assumption to incorporate multiple policies. The key difference is that carryover effects are limited to the previous $l$ time-periods for both policies.
Assumption 12 (Ignorability).
$Y_{it}(\bar{a},\bar{m})\perp(A_{it},M_{it})\mid\bar{Y}_{it-1}(\bar{0},\bar{0}),\qquad\forall\bar{a},\bar{m}$
The multiple policy ignorability assumption simply states that both policies are jointly randomized with respect to the potential outcomes given the potential outcomes absent treatment. Note that this assumption allows for causal pathways between the two policies, so that $M$ ($A$) may lie on the pathway from $A$ ($M$) to $Y$.
Assumption 13 (Constant treatment effects).
$Y_{it}(\bar{a},\bar{m})=Y_{it}(0,0)+\Delta_{t}(\bar{a},\bar{m})$
This assumption again implies that there is no effect heterogeneity on the additive scale (though this could be modified to the multiplicative scale if desired).
Assumption 14 (Positivity).
For some $\epsilon>0$ and $e(y)=P\left\{\bar{A}_{it},\bar{M}_{it}\mid Y_{i,t-1}(\bar{0})=y\right\}$, we have
$P[\min_{\bar{a},\bar{m}}P\left[e\{Y_{i,t-1}(0)\}\geq\epsilon\right]]=1.$
This assumption again simply extends the previous positivity assumption by stating that there is some positive probability of observing any treatment history value for either treatment conditional on a specific value of the previous potential outcome absent any treatments. Under these assumptions, we then obtain the identification result as follows:
$\Delta_{t}$ $=\mathbb{E}[\mathbb{E}[Y_{it}\mid\bar{A}_{it}=\bar{a},\bar{M}_{it}=\bar{m},Y_{i,t-1}(0)]-\mathbb{E}[Y_{it}\mid\bar{A}_{it}=\bar{0},\bar{M}_{it}=\bar{0},Y_{i,t-1}(0)]]$
$=\mathbb{E}[\mathbb{E}(Y_{i,t}\mid\bar{A}_{it}=\bar{a},\bar{M}_{it}=\bar{m},Y_{i,t-1}-\Delta_{t-1}(\bar{A}_{t-1},\bar{M}_{t-1}))$
$-\mathbb{E}(Y_{i,t}\mid\bar{A}_{it}=\bar{0},\bar{M}_{it}=\bar{0},Y_{i,t-1}-\Delta_{t-1}(\bar{A}_{t-1},\bar{M}_{t-1}))].$
Appendix B More details on differential measurement error
In this section, we allow the measurement error to depend on any of the terms in the model, and we highlight the bias obtained in this more general scenario than what was considered in the manuscript. This results in a more general expression of the bias compared to the bias expression seen in the manuscript. Specifically, one can show that our estimate of the treatment effect in the presence of measurement error induced by treatment effect heterogeneity converges to
$\frac{\text{cov}(A_{it},Y_{it})-\frac{\text{cov}(A_{it},Y_{i,t-1}(0)+U_{i,t-1})\text{cov}(Y_{it},Y_{i,t-1}(0)+U_{i,t-1})}{\text{var}(Y_{i,t-1}(0)+U_{i,t-1})}}{\text{var}(A_{it})-\frac{\text{cov}^{2}(A_{it},Y_{i,t-1}(0)+U_{i,t-1})}{\text{var}(Y_{i,t-1}(0)+U_{i,t-1})}}.$
If the measurement error is classical, then any covariance terms involving $U_{i,t-1}$ are zero and therefore the measurement error only impacts the $\text{var}(Y_{i,t-1}(0)+U_{i,t-1})$ term, which simplifies the formula to what was seen in equation (4) of the main manuscript. If, however, the measurement error is differential, the bias can take different magnitudes or directions. Differential measurement error here could occur if $U_{i,t-1}$ is correlated with any of $A_{it},Y_{it},$ or $Y_{i,t-1}(0)$. Therefore it is worth considering in our context whether we expect the bias to be differential, if we have any prior knowledge about the direction of bias, and whether it depends on the source of heterogeneity.
We discussed the impact of correlation between $U_{i,t-1}$ and $Y_{it}$ in the manuscript and the possible bias that can occur in that scenario. Here, we can discuss differential measurement error caused by correlation between $U_{i,t-1}$ and either $A_{it}$ or $Y_{it}(0)$. We believe it is unlikely to have dependence between $A_{it}$ and $U_{i,t-1}$ as units are not likely to be treated based on the magnitude of the individual level treatment effect. If this were to occur, however, it would likely stem from positive dependence between $A_{it}$ and $U_{i,t-1}$, the implications of which would be essentially identical to what was discussed in the manuscript with differential measurement error caused by dependence between $Y_{it}$ and $U_{i,t-1}$. This would lead to treatment effect estimates that are less than what would be seen for the same magnitude of measurement error if it were instead classical.
The last form of differential measurement error to consider is the case where $U_{i,t-1}$ is correlated with $Y_{i,t-1}(0)$. This could occur if units that have larger values of $Y_{it}(0)$ are more likely to benefit from treatment. In the opioid setting, states with higher mortality from opioid use might be more likely to benefit from polices reducing its impact. One thing to note is that dependence between $U_{i,t-1}$ and $Y_{i,t-1}(0)$ makes it very likely to have dependence between $U_{i,t-1}$ and $Y_{it}$ as well given the clear relationship between potential outcomes and observed outcomes. For this reason, many of the issues mentioned in the manuscript for dependence with $Y_{it}$ would also apply in this situation; however, here we focus specifically on the impacts of the dependence between $U_{i,t-1}$ and $Y_{i,t-1}(0)$. Letting $P_{y}(U_{i,t-1})$ denote the projection of $U_{i,t-1}$ onto $Y_{i,t-1}(0)$, the variable we adjust for can be written as
$Y_{i,t-1}(0)+U_{i,t-1}$ $=Y_{i,t-1}(0)+P_{y}(U_{i,t-1})+U_{i,t-1}-P_{y}(U_{i,t-1})$
$=Y_{i,t-1}^{*}(0)+U_{i,t-1}^{*},$
where $Y_{i,t-1}^{*}(0)=kY_{i,t-1}(0)$ for some constant $k$, and $U_{i,t-1}^{*}=U_{i,t-1}-P_{y}(U_{i,t-1})$. Importantly, we have that $\text{Cov}(Y_{i,t-1}^{*}(0),U_{i,t-1}^{*})=0$, which means that differential measurement error with respect to $Y_{i,t-1}(0)$ can be equivalently thought of as measurement error that is not correlated with $Y_{i,t-1}^{\star}(0)$ and the degree of error is driven by the magnitude of $U_{i,t-1}^{*}$. The discussion on classical measurement error therefore holds in this setting; however, importantly, $\text{Var}(U_{i,t-1}^{*})\leq\text{Var}(U_{i,t-1})$. The bias should therefore be smaller in this setting relative to error of the same magnitude that isn’t correlated with $Y_{i,t-1}(0)$.
Appendix C Inference on sample level estimands
In our simulation studies and data analysis, we used our model to estimate quantities other than population average treatment effects. This is particularly useful for comparing our approach to existing estimators such as synthetic controls and DiD estimators, which commonly target average treatment effects on the treated over specific time periods for each unit. For example, one frequently encountered estimand is a sample average treatment effect on the treated. This is defined as,
$SATT_{t}=\frac{1}{n_{t}}\sum_{i\in\mathcal{N}_{t}}Y_{it}(1)-Y_{it}(0),$
where $\mathcal{N}_{t}$ is the set of indices corresponding to the treated units and $n_{t}$ is the number of treated units at time $t$ (note here that we assume $l=0$ and therefore the potential outcome at time $t$ only depends on treatment at time $t$).
In principle it is straightforward to target sample estimands with our model; however, inference requires additional consideration. For more information on distinctions between sample and population estimands, as well as Bayesian inference for them, see Ding and Li (2018) and Li, Ding and Mealli (2023). We consider two natural approaches to inference in this setting that both require calculating the posterior distribution of
$\frac{1}{n_{t}}\sum_{i\in\mathcal{N}_{t}}Y_{it}^{*}(1)-Y_{it}^{*}(0).$
These approaches differ, however, in defining the posterior distribution for $Y_{it}^{*}(a)$. The first approach sets $Y_{it}^{*}(A_{it})=Y_{it}$ and draws $Y_{it}^{*}(1-A_{it})$ from an appropriate distribution (for example, normal or negative binomial) with parameters given by our model. This approach acknowledges that one of the counterfactuals is observed and does not need to be imputed, and then imputes the other from the model. The second approach draws both $Y_{it}^{*}(1)$ and $Y_{it}^{*}(0)$ from the posterior predictive distribution; for example, a normal or negative binomial distribution with parameters given by our model. This approach ignores information from the observed counterfactual and instead imputes both assuming they are independent. The first approach may seem more natural as it uses the observed potential outcome. In our experience, however, this approach works well when the model is correctly specified, but suffers more from moderate amounts of model misspecification relative to the second approach. Many components of the model may be misspecified, including the treatment effect function. If components other than the treatment effect portion of the model are misspecified, we will obtain poor predictions of $Y_{it}^{*}(1-A_{it})$, in turn leading to poor estimates of $SATT_{t}$. The second approach instead uses the difference in two posterior predictive draws, $Y_{it}^{*}(1)-Y_{it}^{*}(0)$: this difference may better estimate the treatment effect, even when individually each of the draws $Y_{it}^{*}(a)$ are biased for the true potential outcome. While both approaches are reliant to some degree on correct model specification, the first approach relies on estimating the full data generating process well, while the second is slightly less reliant on the entire model, but more on the treatment effect specification.
Appendix D Additional information and results for simulation study
D.1 Data generation mechanism for simulation study
Here, we describe in more detail the specifics of the simulation study used in the manuscript. To vary the degree of confounding, we assign treatment to units with higher values of the outcome, and allow treatment times to vary as a function of the outcome as well. We first calculate the outcome for each unit in time period 7, and then used max-min scaling to let all these values fall between 0 and 1. We refer to this value as $\widetilde{Y}_{7}$. To generate start times that better reflects staggered adoption, we randomly sample $20$ numbers between 8 and 14 with replacement and sort them in increasing order. We can refer to this ordered vector of potential times as $\widetilde{T}_{1},\dots,\widetilde{T}_{20}$ where we have that $\widetilde{T}_{1}\leq\widetilde{T}_{2}\leq\dots\leq\widetilde{T}_{20}$. We assign the $j^{th}$ time $\widetilde{T}_{j}$ to be the start time for unit $i$ with probability proportional to $\widetilde{Y}_{7,i}^{c}$, where $c$ is what we refer to as the confounding constant. Once a unit has been assigned a start time, they are removed from the pool of units moving forward and we continue until 20 units have a treatment time. This process ensures that units with higher values of the outcome are more likely to initiate treatment earlier, and the units that are treated have higher outcome values generally. When $c=0$ there is no confounding by $Y_{it}(0)$ and units and times are chosen completely randomly, but as $c$ increases, the degree of confounding increases. To facilitate comparison, we also ensure that units remain treated for all time periods after treatment initiation, though this is not required for our approach.
We also need to generate a treatment effect for each unit in the sample. To ensure that larger states have larger treatment effects, we first work on the rate scale, which is the number of outcomes divided by the population size of each state. We assume the effect is homogeneous with respect to time on the rate scale, but is unique to each state. Formally, we generate new rates $R_{it}^{*}$ as a function of the observed rates $R_{it}$ via
$R_{it}^{*}=R_{it}+A_{it}(\tau+\tau_{i}).$
Throughout, we set $\tau=-0.02$ to fix a small average impact of the policy on rates of prescribing, and we vary $\tau_{i}$ depending on what level of heterogeneity we are considering. We set $\tau_{i}=k\boldsymbol{X}_{i}\boldsymbol{\gamma}$ where $\boldsymbol{X}_{i}$ are time-invariant characteristics, and the elements of $\boldsymbol{\gamma}$ are generated from independent standard normal distributions. We explore $k\in\{0,0.009,0.026\}$, which we denote by no heterogeneity, moderate heterogeneity, and high heterogeneity. The two non-zero values were chosen to ensure that $P(|\tau_{i}|>|\tau|)$ was equal to 0.1 or 0.33, respectively. This implies that in the high heterogeneity setting, approximately one third of the subjects have unit-specific treatment effects that are either of the opposite sign of the average treatment effect, or are at least twice as large. The observed outcomes are set as $Y_{it}^{*}=R_{it}^{*}P_{it}$, where $P_{it}$ is the population size for unit $i$ at time $t$, and outcomes are rounded to the nearest integer to ensure count outcomes.
D.2 Negative binomial simulations on opioid prescribing
Here we evaluate the performance of the proposed autoregressive approach when using negative binomial models. As in the simulation section of the manuscript, we utilize the observed opioid data to perform simulation studies instead of using fully synthetic data. We follow the same simulation structure as in Section 5 of the manuscript that allows for staggered adoption of the policy and confounding by previous time points. The only difference between the data generation in the manuscript and in this section is how the treatment effects are generated. As in the manuscript, a small average treatment effect is incorporated, and we vary the degree of heterogeneity of the treatment effect across units. We choose the degree of heterogeneity in the same manner as in the manuscript, but we generate new rates on the multiplicative scale via:
$\frac{R_{it}^{*}}{R_{it}}=\exp\big{(}A_{it}(\tau+\tau_{i})\big{)},$
so that Assumption 9 is satisfied when there is no heterogeneity ($\tau_{i}=0$ for all $i$). As in the manuscript, the constant treatment effects assumption will be violated in the moderate and high heterogeneity situations. We target the same estimand as in Section 5, given by
$\frac{1}{n_{t}}\sum_{i\in\mathcal{N}_{t}}\sum_{m=0}^{2}\bigg{\{}Y_{i,T_{i}+m}(1)-Y_{i,T_{i}+m}(0)\bigg{\}}.$
which targets the average treatment effect on the treated units in the first three time periods after policy adoption. We compare our approach using negative binomial autoregressive models as described in Section 3.4 of the manuscript with the difference in differences, synthetic control, and two-way fixed effects approaches described in Section 5.
The results can be found in Figures 10 - 13. Overall, the results mirror those seen in Section 5 when using the linear autoregressive models rather than the negative binomial autoregressive models. The improved efficiency leads to large decreases in MSE compared with existing estimators. The synthetic control estimator performs slightly better in terms of MSE in some of the high heterogeneity simulations, though the autoregressive model has the lowest MSE in all other scenarios. The interval coverage for the autoregressive model is near the nominal rate, except in situations with very large degrees of heterogeneity, which echoes the same findings from Section 5 of the manuscript. Lastly, the autoregressive approach has the highest power of all four of the approaches to identify a treatment effect across all simulation scenarios, due to the improved efficiency of the approach.
Figure 10: Distribution of point estimates for simulations on opioid prescribing data with negative binomial autoregressive models.
Figure 11: 95% interval coverage for simulations on opioid prescribing data with negative binomial autoregressive models.
Figure 12: Mean squared error for simulations on opioid prescribing data with negative binomial autoregressive models. MSE values have been capped at 5 for easier comparison.
Figure 13: Power for simulations on opioid prescribing data with negative binomial autoregressive models.
Appendix E Additional information and results for opioid analysis
In this section, we provide some additional information about the data used in Section 6, as well as some additional results that were not shown in the manuscript. Figure 14 shows time trends for the rate of opioid prescriptions by state for each year in the study, with the solid black line representing the nationwide average. We see that in the early years of the study opioid prescriptions were rising, though these began to fall in the second half of the study. Figure 15 shows the distribution of the model parameters corresponding to the treatment effect in the multiple policy setting. The red histogram shows that the effect of must access PDMP laws reduces the rate of opioid prescriptions, the blue histogram shows that medical marijuana has essentially no effect on prescription rates, and the interaction term in green shows that the effect of must access PDMP laws is indeed modified by the presence of medical marijuana laws.
Figure 14: Trend of Opioid Prescribing Rates per 100,000 people, separated by state. The dark line represents the national average.
Figure 15: Posterior distribution of treatment effect parameters in the autoregressive model for the analysis of PDMP and medical marijuana laws jointly.